x/b+c=y/c+a=z/a+bthen prove that.(b-c)x+(c-a)y+(a-b)z=0

Grade:12

1 Answers

Aditya Gupta
4 years ago
let x/(b+c)= y/(c+a)= z/(a+b)= k
so, (b-c)x+(c-a)y+(a-b)z= (b-c)k(b+c) + (c-a)k(c+a) + (a-b)k(a+b)
= k(b^2 – c^2 + c^2 – a^2 + a^2 – b^2)
0 (zero)
kindly approve :)